# Marking Code Without Breaking It: Code Watermarking for Detecting LLM-Generated Code

### v3

@@ Abstract
Identifying LLM-generated code through watermarking poses a challenge in preserving functional correctness. Previous methods rely on the assumption that watermarking high-entropy tokens effectively maintains output quality. Our analysis reveals a fundamental limitation of this assumption: syntax-critical tokens such as keywords often exhibit the highest entropy, making existing approaches vulnerable to logic corruption. We present STONE, a syntax-aware watermarking method that embeds watermarks only in non-syntactic tokens and preserves code integrity. For its rigorous assessment, we also introduce STEM, a comprehensive framework that balances three critical dimensions: correctness, detectability, and imperceptibility. Across Python, C++, and Java, STONE preserves correctness, sustains strong detectability, and achieves balanced performance with minimal overhead. Our implementation is available at https://anonymous.4open.science/r/STONE-watermarking-AB4B/.
@@ 1 Introduction
The rapid development of LLMs has significantly improved the ability to generate human-like text and code (Nam et al., 2024; Kazemitabaar et al., 2023; Liu et al., 2023a; Sun et al., 2024; Guo et al., 2024a; Wang et al., 2024; Wei et al., 2023; Tang et al., 2023; Guo et al., 2024b). These advancements have unlocked new possibilities across various fields, including software development and automated code generation. However, this progress has raised new challenges in tracing the provenance of generated code (Yang et al., 2023; Li et al., 2023). As LLM-generated code becomes increasingly indistinguishable from human-written code, researchers have explored watermarking techniques to address this issue. As illustrated in Figure 1, existing approaches often modify syntax-critical tokens such as keywords, and this modification introduces syntax errors or alters program behavior. In contrast, STONE preserves these tokens and embeds watermarks only into non-syntactic elements, thereby maintaining structural and functional integrity.
@@
LLM watermarking involves embedding specific patterns into the generated output during the generation process to help identify its origin (Sander et al., 2024; Golowich and Moitra, 2024; Hu and Huang, 2024; Giboulot and Furon, 2024; Li et al., 2024b; Pang et al., 2024; Zhou et al., 2024; Fu et al., 2024; Panaitescu-Liess et al., 2025). These patterns remain imperceptible to humans and detectable through algorithmic analysis, thereby promoting greater transparency and accountability in AI-generated content (Guo et al., 2024c; Hou et al., 2024). Among various approaches, the green and red list-based watermarking is the most widely adopted (Kirchenbauer et al., 2023; Zhao et al., 2023; Lee et al., 2024; Lu et al., 2024; Guo et al., 2024c). These approaches divide token candidates into two lists (green and red) and biases the selection process toward tokens in the green list. Detection involves measuring the proportion of green list tokens in the generated output. While effective for natural language tasks, this approach does not easily apply to code generation, where structural and functional correctness are crucial (Lee et al., 2024; Hoang et al., 2024). Changes to syntax-critical tokens can cause compilation errors or alter program behavior.
We propose STONE (Syntax TOkeN preserving codE watermarking), a method that selectively embeds watermarks in non-syntax tokens. By excluding tokens that are critical to code execution, STONE reduces the risk of functional degradation while embedding robust, detectable patterns.
Additionally, we introduce STEM (Summative Test framework for Evaluating code-waterMarking). Existing studies on code watermarking rely on separate evaluation criteria, using different metrics and measurement protocols across methods, hindering fair and consistent comparison of performance. STEM integrates correctness, detectability, and imperceptibility into a unified framework, providing a fair basis for evaluating overall performance while allowing task-specific weighting when needed. We employ STEM to evaluate STONE and compare it with state-of-the-art watermarking methods. Experiments across Python, C++, and Java show that STONE preserves functional correctness, achieves balanced performance across all evaluation criteria, and consistently outperforms prior methods in efficiency.
@@ 2 Motivation and Preliminary Analysis
The recent code watermarking method SWEET (Lee et al., 2024) accounts for structural and syntactic constraints by embedding watermarks only in high-entropy tokens, thereby minimizing their impact on code quality. In our preliminary analysis, we group code tokens into several categories and compare their average entropy values to see if high-entropy tokens have a smaller impact on code quality. We consider two popular benchmarks, MBPP+ and HumanEval+, for the Python code generation and completion. We classify each Python code into one of five categories (keywords, whitespace, types, delimiters, and operators) where any remaining tokens are put into the etc. category. A detailed explanation is provided in Appendix A.
We calculate the entropy values of tokens across different categories and compare the average entropy values for each category. Following Lee et al. (2024), we calculate the Shannon entropy. Given a token sequence $y=(y_{0},y_{1},\ldots,y_{N})$, the entropy at each generation step $t$ is computed based on the predicted probability distribution of the next token. The entropy $H_{t}$ is defined as:
@@
where $H_{t}$ is the entropy value at generation step $t$, $V$ is the vocabulary of the language model, $y_{<t}$ is the sequence of tokens generated before step $t$, and $P(y_{t}=v_{i}\mid y_{<t})$ is the probability assigned by the model to token $v_{i}$ as the next token. The entropy $H_{t}$ measures the uncertainty of the next token prediction. A high entropy value indicates that the model is uncertain and many tokens have comparable probabilities. Conversely, low entropy implies that one token has a dominant probability.
Figure 2 presents bar charts displaying the mean entropy values for each token category. Our analysis shows that the average entropy of keyword tokens is the highest among all categories. In Python code, keywords are used in various contexts and positions. For example, control flow keywords like if, for, and while can appear in different parts of the code with highly diverse usage patterns. In contrast, operators and delimiters follow more predictable patterns due to the syntactic structure of the code. For instance, operators such as +, -, and * are used between operands, while delimiters like ( and ) appear in specific patterns, such as in function or class definitions. Additionally, type tokens are mainly used in limited contexts, such as variable declarations or type hinting, making their occurrence and patterns relatively predictable. Whitespace tokens, following the structure of the code, are even more predictable in their placement.
Since keywords have high entropy, embedding watermarks in high-entropy tokens, as in SWEET, carries the risk of placing watermarks on keywords. Keywords are essential to the syntax and logic of the code, thus watermarking them could negatively impact the quality and functionality of the code. In comparison, the etc category, while less critical to the functionality of the code, has relatively high entropy. Therefore, embedding watermarks in etc tokens (i.e., non-syntactic tokens) may be a more suitable approach for preserving code quality.
@@ 3.1 Syntax-Aware Watermarking: STONE
STONE is a code watermarking method specifically designed to preserve the quality of generated code. This section describes how STONE performs watermark insertion and detection. We describe the insertion process, where watermarks are embedded only in non-syntactic tokens while keeping critical elements such as keywords, operators, and delimiters unchanged to preserve functionality and structure. Detection is performed by analyzing the distribution of green tokens among non-syntax elements.
@@
We define the keywords, whitespace, types, delimiters, and operators as functionally important syntactic elements and embed watermarks only in tokens that do not belong to these categories. Algorithm 1 describes how STONE embeds watermarks during code generation. Given a prompt $\mathbf{x}$ and $\mathbf{y}_{:t}$, the tokens generated up to the previous time step, the language model $f_{\text{LM}}$ computes the initial probability distribution $\mathbf{p}_{t}$ for the token to be generated at the current time step $t$. STONE samples a candidate token $\tilde{y}_{t}$ based on this initial probability distribution. If the sampled token is not in the syntax element set, STONE divides the vocabulary $\mathcal{V}$ into a green/red list according to the green token ratio $\gamma$. It then increases the likelihood of sampling tokens from the green list by adding a constant $\delta$ to their logit values. Using the adjusted logits, STONE updates the probability distribution and samples the final token $y_{t}$ for time step $t$.
@@
1:Prompt $\mathbf{x}$, previously generated tokens $\mathbf{y}_{:t}=(y_{0},\dots,y_{t-1})$, vocabulary set $\mathcal{V}$, language model $f_{\text{LM}}$, syntax element set $S$, green list ratio $\gamma\in(0,1)$, logit adjustment factor $\delta>0$
2:for $t=0,1,2,\dots$ do
3:  Compute the logit vector: $\mathbf{l}_{t}=f_{\text{LM}}(\mathbf{x},\mathbf{y}_{:t})$
@@
$p_{t,i}=\frac{e^{l_{t}[i]}}{\sum_{j=1}^{|\mathcal{V}|}e^{l_{t}[j]}},\quad\text{for }i\in\{0,\dots,|\mathcal{V}|-1\}$
@@
7:   Compute a hash of token $y_{t-1}$ to use as a seed
8:   Split $\mathcal{V}$ into $\mathcal{G}_{t}$ and $\mathcal{R}_{t}$:
$\mathcal{G}_{t}\quad(\text{green list of size }\gamma|\mathcal{V}|)$
$\mathcal{R}_{t}\quad(\text{red list of size }(1-\gamma)|\mathcal{V}|)$
9:   Add $\delta$ to the logits of tokens in $\mathcal{G}_{t}$:
$l_{t}[i]\leftarrow l_{t}[i]+\delta,\quad\text{for }i\in\mathcal{G}_{t}$
@@
$p^{\prime}_{t,i}=\frac{e^{l_{t}[i]}}{\sum_{j=1}^{|\mathcal{V}|}e^{l_{t}[j]}},\quad\text{for }i\in\{0,\dots,|\mathcal{V}|-1\}$
@@
1:Token sequence to be tested $\mathbf{X}=(X_{0},\dots,X_{N-1})$, syntax element set $S$, green list ratio $\gamma\in(0,1)$, $z$-score threshold $z_{\text{threshold}}>0$
2:Set $N^{E}\leftarrow 0$, $N^{E}_{G}\leftarrow 0$
3:for $t=0,1,\dots,N-1$ do
4:  if $X_{t}\notin S$ then $\triangleright$ Non-syntax element detected
@@
6:   Compute a hash of token $X_{t-1}$ to use as a seed
7:   Split $\mathcal{V}$ into $\mathcal{G}_{t}$ and $\mathcal{R}_{t}$:
$\mathcal{G}_{t}\quad(\text{green list of size }\gamma|\mathcal{V}|)$
$\mathcal{R}_{t}\quad(\text{red list of size }(1-\gamma)|\mathcal{V}|)$
@@
9:     $N^{E}_{G}\leftarrow N^{E}_{G}+1$ $\triangleright$ Token matches the green list
@@
13:Compute the $z$-score: $z=\frac{N^{E}_{G}-\gamma N^{E}}{\sqrt{\gamma(1-\gamma)N^{E}}}$
@@
15:  return True $\triangleright$ The sequence $\mathbf{X}$ is watermarked
@@
17:  return False $\triangleright$ The sequence $\mathbf{X}$ is not watermarked
@@
3.2 Unified Evaluation Framework: STEM
Prior studies on code watermarking have typically evaluated functional correctness, detectability, and imperceptibility as separate criteria (Lee et al., 2024; Yang et al., 2024; Li et al., 2024a; Guan et al., 2024). However, evaluating each criterion in isolation makes it difficult to capture the trade-offs between them and to identify methods with balanced performance. For example, a watermarking technique may enhance detectability at the expense of correctness, resulting in incomplete or conflicting assessments. This work addresses this limitation by introducing STEM, a balanced evaluation framework that integrates correctness, detectability, and imperceptibility into a unified score. STEM combines the three criteria using weights $\alpha$, $\beta$, and $\zeta$ satisfying $\alpha+\beta+\zeta=1$, enabling direct comparison of methods in terms of overall balance rather than isolated metrics.
@@ 3.2 Unified Evaluation Framework: STEM
Maintaining correctness is crucial for preventing watermarking from compromising code functionality. Let $C_{wm}$ denote the set of watermarked code samples. We measure functionality preservation using the unbiased estimator of pass@k Chen et al. (2021):
@@
where $n$ and $c$ denote the total and correct generations per task, respectively. A higher score indicates that watermark insertion does not hinder program execution.
@@
Detectability measures how effectively a watermark can be identified in generated code. We assess detectability using a $z$-score that quantifies the proportion of green-list tokens and compute AUROC between human-written and watermarked corpora.
@@
where $|X_{wm}|_{G}$ is the number of green tokens in $X_{wm}$ and $\gamma$ is the expected green-token ratio in non-watermarked code. Higher $z$ values indicate stronger evidence that the code is watermarked.
Given a corpus of human-written code $C_{H}=\{X^{(i)}_{H}\}$ and watermarked code $C_{wm}=\{X^{(j)}_{wm}\}$, we compute the true- and false-positive rates at threshold $\tau$:
@@
$\mathrm{Detectability}(C_{wm},C_{H})=\int_{0}^{1}\text{TPR}(\text{FPR})\,d(\text{FPR}),$ (7)
where higher AUROC values indicate stronger detectability between watermarked and human-written code.
@@
Watermarking should preserve token probability distributions consistent with normal model outputs (Huang and Wan, 2025). In code watermarking, Sun et al. (2023) assessed imperceptibility through post-hoc detection combining human and automated evaluations, which does not quantify distributional distortion. Following perplexity-based approaches in text watermarking (Kirchenbauer et al., 2023) and studies including code generation (Magnusson et al., 2024), we define imperceptibility as the deviation in token probability distributions introduced by watermarking, measured through perplexity shifts computed with code LLMs.
Let $C_{wm}$ and $C$ denote the corpora of watermarked and non-watermarked code, respectively. Perplexity is computed as
$\operatorname{PPL}(C_{wm})=\begin{aligned} &\frac{1}{\lvert C_{wm}\rvert}\sum_{j=1}^{\lvert C_{wm}\rvert}\exp\!\bigl(-A_{j}\bigr),\\ &A_{j}=\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\log P\!\left(y_{i}^{(j)}\mid y_{<i}^{(j)}\right).\end{aligned}$ (8)
Higher values indicate that watermark insertion produces little distributional deviation from normal code generation.

### v4

@@ Abstract
Identifying LLM-generated code through watermarking poses a challenge in preserving functional correctness. Previous methods rely on the assumption that watermarking high-entropy tokens effectively maintains output quality. Our analysis reveals a fundamental limitation of this assumption: syntax-critical tokens such as keywords often exhibit the highest entropy, making existing approaches vulnerable to logic corruption. We present STONE, a syntax-aware watermarking method that embeds watermarks only in non-syntactic tokens and preserves code integrity. For rigorous evaluation, we also introduce STEM, a comprehensive metric that balances three critical dimensions: correctness, detectability, and imperceptibility. Across Python, C++, and Java, STONE preserves correctness, sustains strong detectability, and achieves balanced performance with minimal computational overhead. Our implementation is available at https://github.com/inistory/STONE-watermarking.
@@ 1 Introduction
The rapid development of LLMs has significantly improved their ability to generate human-like code  (Zan et al., 2022; Tang et al., 2023; Wei et al., 2023). These advancements have unlocked new possibilities, including software development and automated code generation (Nam et al., 2024; Wang et al., 2024; Guo et al., 2024a). However, this progress has raised new challenges in tracing the provenance of generated code (Yang et al., 2023; Li et al., 2023). As LLM-generated code becomes increasingly indistinguishable from human-written code, researchers have explored watermarking techniques to address this issue. As illustrated in Figure 1, existing approaches often modify syntax-critical tokens, which can introduce syntax errors or alter program behavior. In contrast, STONE preserves these tokens and embeds watermarks only into non-syntactic elements, thereby maintaining structural and functional integrity.
@@
LLM watermarking embeds specific patterns into the generated output during the generation process to help identify its origin  (Sander et al., 2024; Golowich and Moitra, 2024; Hu and Huang, 2024) These patterns remain imperceptible to humans yet detectable through algorithmic analysis, thereby promoting greater transparency and accountability in AI-generated content (Guo et al., 2024b; Hou et al., 2024). Among various approaches, the green and red list-based watermarking is the most widely adopted (Kirchenbauer et al., 2023; Zhao et al., 2023; Lee et al., 2024; Lu et al., 2024). These approaches divide token candidates into two lists (green and red) and bias the selection process toward tokens in the green list. Detection involves measuring the proportion of green list tokens in the generated output. While effective for natural language tasks, this approach does not easily apply to code generation, where structural and functional correctness are crucial (Lee et al., 2024; Hoang et al., 2024). Changes to syntax-critical tokens can cause compilation errors or alter program behavior.
We propose STONE (Syntax TOkeN preserving codE watermarking), a method that selectively embeds watermarks in non-syntax tokens. By excluding tokens that are critical to code execution, STONE reduces the risk of functional degradation while embedding robust, detectable patterns. Additionally, we introduce STEM (Summative Test metric for Evaluating code-waterMarking). Existing studies on code watermarking rely on separate evaluation criteria, using different metrics and measurement protocols across methods, hindering fair and consistent comparison of performance. STEM integrates correctness, detectability, and imperceptibility into a unified metric, providing a fair basis for evaluating overall performance while allowing task-specific weighting. We employ STEM to evaluate STONE and compare it with state-of-the-art watermarking methods. Experiments across Python, C++, and Java show that STONE preserves functional correctness, achieves balanced performance across all evaluation criteria, and consistently outperforms prior methods in efficiency.
@@ 2 Motivation and Preliminary Analysis
The recent code watermarking method SWEET (Lee et al., 2024) embeds watermarks in high-entropy tokens. This design is motivated by the intuition that modifying tokens with higher uncertainty is less likely to affect the perceived quality of the generated code. In our preliminary analysis, we examine whether token entropy indeed varies systematically across syntactic categories and how such variation relates to potential impacts on code quality.
We consider two popular benchmarks, MBPP+ and HumanEval+, for Python code generation and completion. Each Python token is classified into one of five syntactic categories (keywords, whitespace, types, delimiters, and operators), with all remaining tokens grouped into an etc category. Details of the categorization procedure are provided in Appendix C.
We compute token entropy for each category using Shannon entropy, following Lee et al. (2024). Given a token sequence $y=(y_{0},y_{1},\ldots,y_{N})$, the entropy at generation step $t$ is defined as:
@@
where $V$ denotes the vocabulary and $y_{<t}$ the previously generated tokens. This formulation captures the uncertainty of the model’s next-token prediction at each generation step.
Figure 2 shows the mean entropy values for each token category. We observe that keywords exhibit the highest average entropy among all categories. This can be attributed to their diverse roles across control flow, structural definitions, and other syntactic constructs in Python. In contrast, operators, delimiters, types, and whitespace follow more regular and constrained syntactic patterns, resulting in comparatively lower entropy.
Although keywords have high entropy, watermarking them risks altering tokens that are essential to code syntax and semantics. Even small perturbations to such tokens may lead to functional errors or noticeable degradation in code quality. In comparison, the etc category contains tokens that are less critical to core syntactic structure while still exhibiting relatively high entropy. This observation suggests that embedding watermarks in non-syntactic tokens may preserve code quality better than approaches that rely solely on high-entropy tokens.
@@ 3.1 Syntax-Aware Watermarking: STONE
STONE is a code watermarking method designed to preserve the quality of generated code. This section introduces the watermark insertion and detection procedures of STONE.
@@
We treat keywords, whitespace, types, delimiters, and operators as syntactic elements and embed watermarks only in non-syntactic tokens. Algorithm 1 describes how STONE embeds watermarks during code generation. Given a prompt $\mathbf{x}$ and previously generated tokens $\mathbf{y}_{<t}$, the language model $f_{\text{LM}}$ computes the initial probability distribution $\mathbf{p}_{t}$ at time step $t$. STONE first samples a candidate token $\tilde{y}_{t}$ from the initial probability distribution. If the sampled token is not in the syntax element set, STONE then divides the vocabulary $\mathcal{V}$ into a green and red list according to the green token ratio $\gamma$. It then increases the likelihood of sampling green-list tokens by adding a constant $\delta$ to their logit values. Using the adjusted logits, STONE updates the probability distribution and samples the final token $y_{t}$ for time step $t$.
@@
1:Prompt $\mathbf{x}$, start token $y_{0}$, last generated tokens $\mathbf{y}_{<t}=(y_{0},\dots,y_{t-1})$, vocabulary set $\mathcal{V}$, language model $f_{\text{LM}}$, syntax element set $S$, green list ratio $\gamma\in(0,1)$, logit adjustment factor $\delta>0$
2:for $t=1,2,3,\dots$ do
3:  Compute the logit vector:
$\mathbf{l}_{t}=f_{\text{LM}}(\mathbf{x},\mathbf{y}_{<t})$
@@
$p_{t,i}=\frac{e^{\mathbf{l}_{t}[i]}}{\sum_{j=1}^{|\mathcal{V}|}e^{\mathbf{l}_{t}[j]}},\quad\text{for }i\in\{1,\dots,|\mathcal{V}|\}$
@@
7:   Compute a hash of token $y_{t-1}$ as a seed
8:   Split vocabulary indices into $\mathcal{G}_{t}$ and $\mathcal{R}_{t}$
9:   Increase logits for indices in $\mathcal{G}_{t}$ by $\delta$:
$\mathbf{l}_{t}[i]\leftarrow\mathbf{l}_{t}[i]+\delta,\quad\text{for }i\in\mathcal{G}_{t}$
@@
$p^{\prime}_{t,i}=\frac{e^{\mathbf{l}_{t}[i]}}{\sum_{j=1}^{|\mathcal{V}|}e^{\mathbf{l}_{t}[j]}},\quad\text{for }i\in\{1,\dots,|\mathcal{V}|\}$
@@
1:Token sequence $\mathbf{X}=(X_{0},\dots,X_{N-1})$, syntax element set $S$, green list ratio $\gamma\in(0,1)$, $z$-score threshold $z_{\text{threshold}}>0$
2:Initialize $N^{E}\leftarrow 0$, $N^{E}_{G}\leftarrow 0$
3:for $t=1,2,\dots,N-1$ do
4:  if $X_{t}\notin S$ then
@@
6:   Compute a hash of token $X_{t-1}$ as a seed
7:   Split vocabulary indices into $\mathcal{G}_{t}$ and $\mathcal{R}_{t}$
@@
9:     $N^{E}_{G}\leftarrow N^{E}_{G}+1$
@@
13:Compute the $z$-score:
$z=\frac{N^{E}_{G}-\gamma N^{E}}{\sqrt{\gamma(1-\gamma)N^{E}}}$
@@
15:  return True $\triangleright$ Sequence $\mathbf{X}$ is watermarked
@@
17:  return False $\triangleright$ Sequence $\mathbf{X}$ is not watermarked
@@
3.2 Unified Evaluation Metric: STEM
Prior studies on code watermarking typically assess functional correctness, detectability, and imperceptibility independently (Lee et al., 2024; Yang et al., 2024; Li et al., 2024; Guan et al., 2024). They adopt diverse definitions and evaluation protocols, which prevents consistent and fair comparison across methods. For instance, a technique may improve detectability at the cost of correctness, leading to inconsistent evaluations. STEM addresses this issue by unifying the three criteria into a single weighted metric $(\alpha,\beta,\zeta)$, where $\alpha+\beta+\zeta=1$, enabling fair and direct comparison across methods under configurable trade-offs.
@@ 3.2 Unified Evaluation Metric: STEM
Correctness is crucial for preventing watermarking from compromising code functionality. Let $C_{wm}$ denote the set of watermarked code samples. We measure functionality preservation using the unbiased estimator of pass@k Chen et al. (2021), which is standard for evaluating code generation correctness:
@@
where $n$ corresponds to the total number of generated solutions for a task, and $c$ denotes the subset of those solutions that pass all test cases. This estimator measures the probability that at least one of the $k$ generated solutions is functionally correct. In the context of watermarking, it captures whether watermark insertion introduces execution failures or alters program semantics. A higher correctness score indicates that watermark insertion does not hinder program execution.
@@
Detectability measures how effectively a watermark can be identified in generated code. We assess detectability using a $z$-score that quantifies the proportion of green-list tokens, and aggregate scores by computing the AUROC between human-written and watermarked corpora.
@@
where $|X_{wm}|_{G}$ is the number of green tokens in $X_{wm}$ and $\gamma$ is the expected green-token ratio in non-watermarked code. Higher $z$ values indicate stronger evidence that the code is watermarked. The $z$-score serves as a continuous detection statistic, enabling threshold-based discrimination between watermarked and non-watermarked sequences. By varying the threshold $\tau$, different operating points in terms of true and false positive rates are obtained, forming the ROC curve.
Given a corpus of human-written code $C_{H}=\{X^{(i)}_{H}\}$ and watermarked code $C_{wm}=\{X^{(j)}_{wm}\}$, the true positive rate (TPR) and false positive rate (FPR) at threshold $\tau$ are defined as follows:
@@
$\mathrm{Detectability}(C_{wm},C_{H})=\\ \int_{0}^{1}\text{TPR}(\text{FPR})\,d(\text{FPR}),$ (7)
where higher AUROC values indicate stronger separability between watermarked and human-written code.
@@
Imperceptibility concerns the preservation of token probability distributions consistent with normal model outputs (Huang and Wan, 2025). Since code LLMs follow stable and highly structured token distributions, distortions introduced by watermarking can form statistical signals that are exploitable by adversaries without access to the secret key. Accordingly, following perplexity-based approaches in text watermarking (Kirchenbauer et al., 2023) and studies on code generation (Magnusson et al., 2024), we define imperceptibility as the deviation in token probability distributions introduced by watermarking, measured through perplexity shifts computed with code LLMs.
Let $C_{wm}$ and $C$ denote the corpora of watermarked and non-watermarked code, respectively. Perplexity is computed over the watermarked corpus ${C}_{wm}=\{X_{wm}^{(1)},\dots,X_{wm}^{(J)}\}$ as:
$\operatorname{PPL}(C_{wm})=\\ \frac{1}{\lvert C_{wm}\rvert}\sum_{j=1}^{\lvert C_{wm}\rvert}\exp\!\Bigl(-\mathbb{E}_{i}\!\left[\log P\!\left(y_{i}^{(j)}\mid y_{<i}^{(j)}\right)\right]\Bigr),$ (8)
where $P(y_{i}^{(j)}|y_{<i}^{(j)})$ represents the probability of token $y_{i}^{(j)}$ given its preceding context. An effectively concealed watermark minimizes perplexity shifts, ensuring imperceptibility. We define the imperceptibility metric as:
$\mathrm{Imperceptibility}(C_{wm},C)=\\ 1-\frac{\bigl|\operatorname{PPL}(C_{wm})-\operatorname{PPL}(C)\bigr|}{\operatorname{PPL}(C)}.$ (9)
This metric quantifies the perplexity-based difference between LLM-generated code with watermark ($C_{wm}$) and non-watermarked code ($C$). It computes the relative difference by normalizing the absolute gap by the perplexity of $C$, providing a scale-invariant measure of distributional shift. Higher values indicate minimal impact on token distributions and thus stronger imperceptibility, while lower or negative values reflect greater disruption due to watermark insertion.